# Bilayer Cuprate Antiferromagnets Enable Programmable Cavity Optomagnonics

### v1

@@ IV Spectroscopic Response
Fig. 3 shows the transmission versus magnetic field with the probe locked to the $\alpha$ mode, $f_{\mathrm{probe}}=f_{\alpha}(B)$, for fixed cavity frequencies $f_{c}$. In this configuration, the $\alpha$ branch contributes a purely real (dissipative) self-energy, $\Sigma_{\alpha}(B)=\frac{|g_{\alpha}(B)|^{2}}{\gamma_{\alpha}/2}$, which typically satisfies $\Sigma_{\alpha}>\kappa_{c}$ and therefore sets the effective linewidth. The $\beta$ branch, pinned near $f_{\beta}\simeq 216.98~\mathrm{GHz}$, is far detuned: $\Re[\Sigma_{\beta}]$ adds a small dissipative load and $\Im[\Sigma_{\beta}]$ a dispersive shift; numerical control study setting $g_{\beta}\!\to\!0$ confirms that its influence is negligible at low fields (i.e., away from the $\alpha$$\beta$ degeneracy). Since $|g_{\alpha}|\propto\sqrt{f_{c}/B}$, higher cavity frequencies strengthen $\Sigma_{\alpha}$, broadening the linewidth and reducing the transmission amplitude. This cavity-frequency control of dissipation enables broadband magnon spectroscopy, tunable linewidths, and reconfigurable magnon–photon interfaces. Finite $\mathcal{G}_{\beta}$ produces a clear Fano-type interference at the $\alpha$$\beta$ degeneracy (where $f_{\mathrm{probe}}=f_{\alpha}=f_{\beta}$ and, in our model, $\mathcal{G}_{\alpha}\!\approx\!\mathcal{G}_{\beta}$). As $f_{c}$ increases, the interference visibility grows while the field width changes only weakly, consistent with the Schrieffer–Wolff (Eq.22). Direct microwave access to the $\beta$ mode near $f_{\beta}\simeq 216.98~\mathrm{GHz}$ is impractical; this can be addressed by (i) lowering $\alpha_{D}$ to shift the crossing below $100~\mathrm{GHz}$, or (ii) exploiting coherent $\alpha$$\beta$ coupling for state transfer—mapping states from the field-tunable $\alpha$ mode onto the robust $\beta$ mode for optical retrieval. Increasing $f_{c}$ thus strengthens $\alpha$$\beta$ hybridization, a practical knob for reconfigurable filtering and microwave–optical transduction (with ultimate efficiency set by the engineered optical interface) Liu et al. (2025); Gollwitzer et al. (2021). In the bad-cavity regime [Fig. 3(b)], the qualitative response is preserved, with only the transmission level enhanced, while the linewidth remains unchanged as it is set by $\Sigma_{\alpha}$.
@@
Fig. 4 shows the transmission with the probe locked to the cavity $(f_{\mathrm{probe}}=f_{c})$ as a function of magnetic field $B$. On resonance $(f_{\alpha}(B)=f_{c})$ the $\alpha$-induced loading is strong: the on-resonance self-energy $\Re\Sigma_{\alpha}=|g_{\alpha}|^{2}/(\gamma_{\alpha}/2)\propto f_{c}/B$ suppresses $|S_{21}|^{2}$ well below the bare-cavity baseline $4\kappa_{1}\kappa_{2}/\kappa_{c}^{2}$. As $B$ increases away from resonance, the coupling weakens and the detuning grows, so the loading collapses and $|S_{21}|^{2}$ recovers toward the bare value. In the bad-cavity regime $(\kappa_{c}\gg\gamma_{\alpha})$, this presents as a sharp, field-tunable notch at $f_{\alpha}(B)=f_{c}$ with a magnon-limited linewidth $\simeq\gamma_{\alpha}$, tending toward $\kappa_{c}$ only in the weak-loading limit. At fixed $B$, increasing $f_{c}$ amplifies the $\alpha$-feature $\big(|\mathcal{G}_{\alpha}|\propto\sqrt{f_{c}}\big)$ and shifts the recovery to higher fields. As $f_{c}\to f_{\beta}$, residual $\beta$-mode dressing adds a faint dispersive background with amplitude scaling $\propto|g_{\beta}|^{2}/\big[(\omega_{\beta}-\omega_{c})^{2}+(\kappa_{c}/2)^{2}\big]$. Collectively, these trends yield a compact, field-programmable microwave notch/transparency element, providing a practical spectroscopic handle on the $\alpha$-mode dispersion, damping, and magnetic filling.
@@
Fig. 5 shows the transmission versus probe frequency at fixed $f_{c}=12~\mathrm{GHz}$ while the field sets $\Delta_{\alpha p}\!\equiv\!f_{c}-f_{\alpha}(B)\in\{0,+5,-5\}\,\mathrm{GHz}$. At $\Delta_{\alpha p}=0$ a symmetric normal-mode doublet appears with splitting $2g_{\alpha p}\!\approx\!6.10~\mathrm{GHz}$ (peaks at $8.950$ and $15.047$ GHz); a fit gives $C=4g_{\alpha p}^{2}/(\kappa_{c}\gamma_{\alpha})\approx 7.6\times 10^{1}$ and satisfies the resolvable-splitting criterion $g_{\alpha p}>\lvert\kappa_{c}-\gamma_{\alpha}\rvert/4$, with polariton linewidths $(\kappa_{c}+\gamma_{\alpha})/2\simeq 1.3~\mathrm{GHz}$. For $|\Delta_{\alpha p}|>0$ the $\alpha$ mode induces a dispersive pull $f_{\mathrm{eff}}\simeq f_{c}+\mathrm{Im}\,\Sigma_{\alpha}(f_{c})$, shifting the cavity-like peak up (down) for $\Delta_{\alpha p}>0$ $(<0)$ and producing a Fano-like feature near $f_{\alpha}$ whose asymmetry reverses with the sign of $\Delta_{\alpha p}$. A faint, ultranarrow line at $f_{\beta}$ arises from a virtual-photon pathway via the off-resonant cavity, with power $\propto|g_{\beta}|^{2}/\big[(f_{\beta}-f_{c})^{2}+(\kappa_{c}/2)^{2}\big]$. Together, these effects yield a field-programmable microwave element (band-splitter, notch, or dispersive shifter) and enable single-scan extraction of $g_{\alpha p}$, $\kappa_{c}$, and $\gamma_{\alpha}$, demonstrating GHz-readout of a sub-THz magnon without retuning; in the good-cavity limit the polariton peaks narrow markedly, supporting high-$Q$ band splitting, filtering, and precision spectroscopy Wang et al. (2018); Boventer et al. (2020); Noh et al. (2023).

### v2

@@ IV Spectroscopic Response
Fig. 3 shows the transmission versus magnetic field with the probe locked to the $\alpha$ mode, $f_{\mathrm{probe}}=f_{\alpha}(B)$, for fixed cavity frequencies $f_{c}$. In this configuration, the $\alpha$ branch contributes a purely real (dissipative) self-energy, $\Sigma_{\alpha}(B)=\frac{|g_{\alpha}(B)|^{2}}{\gamma_{\alpha}/2}$, which typically satisfies $\Sigma_{\alpha}>\kappa_{c}$ and therefore sets the effective linewidth. The $\beta$ branch, pinned near $f_{\beta}\simeq 216.98~\mathrm{GHz}$, is far detuned: $\Re[\Sigma_{\beta}]$ adds a small dissipative load and $\Im[\Sigma_{\beta}]$ a dispersive shift; numerical control study setting $g_{\beta}\!\to\!0$ confirms that its influence is negligible at low fields (i.e., away from the $\alpha$$\beta$ degeneracy). Since $|g_{\alpha}|\propto\sqrt{f_{c}/B}$, higher cavity frequencies strengthen $\Sigma_{\alpha}$, broadening the linewidth and reducing the transmission amplitude. This cavity-frequency control of dissipation enables broadband magnon spectroscopy, tunable linewidths, and reconfigurable magnon–photon interfaces. Finite $\mathcal{G}_{\beta}$ produces a clear Fano-type interference at the $\alpha$$\beta$ degeneracy (where $f_{\mathrm{probe}}=f_{\alpha}=f_{\beta}$ and, in our model, $\mathcal{G}_{\alpha}\!\approx\!\mathcal{G}_{\beta}$). As $f_{c}$ increases, the interference visibility grows while the field width changes only weakly, consistent with the Schrieffer–Wolff (Eq.22). Direct microwave access to the $\beta$ mode near $f_{\beta}\simeq 216.98~\mathrm{GHz}$ is impractical; this can be addressed by (i) lowering $\alpha_{D}$ to shift the crossing below $100~\mathrm{GHz}$, or (ii) exploiting coherent $\alpha$$\beta$ coupling for state transfer—mapping states from the field-tunable $\alpha$ mode onto the robust $\beta$ mode for optical retrieval. Increasing $f_{c}$ thus strengthens $\alpha$$\beta$ hybridization, a practical knob for reconfigurable filtering and microwave–optical transduction (with ultimate efficiency set by the engineered optical interface) [15, 97]. In the bad-cavity regime [Fig. 3(b)], the qualitative response is preserved, with only the transmission level enhanced, while the linewidth remains unchanged as it is set by $\Sigma_{\alpha}$.
@@
Fig. 5 shows the transmission versus probe frequency at fixed $f_{c}=12~\mathrm{GHz}$ while the field sets $\Delta_{\alpha p}\!\equiv\!f_{c}-f_{\alpha}(B)\in\{0,+5,-5\}\,\mathrm{GHz}$. At $\Delta_{\alpha p}=0$ a symmetric normal-mode doublet appears with splitting $2g_{\alpha p}\!\approx\!6.10~\mathrm{GHz}$ (peaks at $8.950$ and $15.047$ GHz); a fit gives $C=4g_{\alpha p}^{2}/(\kappa_{c}\gamma_{\alpha})\approx 7.6\times 10^{1}$ and satisfies the resolvable-splitting criterion $g_{\alpha p}>\lvert\kappa_{c}-\gamma_{\alpha}\rvert/4$, with polariton linewidths $(\kappa_{c}+\gamma_{\alpha})/2\simeq 1.3~\mathrm{GHz}$. For $|\Delta_{\alpha p}|>0$ the $\alpha$ mode induces a dispersive pull $f_{\mathrm{eff}}\simeq f_{c}+\mathrm{Im}\,\Sigma_{\alpha}(f_{c})$, shifting the cavity-like peak up (down) for $\Delta_{\alpha p}>0$ $(<0)$ and producing a Fano-like feature near $f_{\alpha}$ whose asymmetry reverses with the sign of $\Delta_{\alpha p}$. A faint, ultranarrow line at $f_{\beta}$ arises from a virtual-photon pathway via the off-resonant cavity, with power $\propto|g_{\beta}|^{2}/\big[(f_{\beta}-f_{c})^{2}+(\kappa_{c}/2)^{2}\big]$. Together, these effects yield a field-programmable microwave element (band-splitter, notch, or dispersive shifter) and enable single-scan extraction of $g_{\alpha p}$, $\kappa_{c}$, and $\gamma_{\alpha}$, demonstrating GHz-readout of a sub-THz magnon without retuning; in the good-cavity limit the polariton peaks narrow markedly, supporting high-$Q$ band splitting, filtering, and precision spectroscopy [98, 99, 100, 101].